### v5

@@ Abstract
We describe a new dependent-rounding algorithmic framework for bipartite graphs. Given a fractional assignment $\vec{x}$ of values to edges of graph $G=(U\cup V,E)$, the algorithms return an integral solution $\vec{X}$ such that each right-node $v\in V$ has at most one neighboring edge $f$ with $X_{f}=1$, and where the variables $X_{e}$ also satisfy broad nonpositive-correlation properties. In particular, for any edges $e_{1},e_{2}$ sharing a left-node $u\in U$, the variables $X_{e_{1}},X_{e_{2}}$ have strong negative-correlation properties, i.e. the expectation of $X_{e_{1}}X_{e_{2}}$ is significantly below $x_{e_{1}}x_{e_{2}}$.
@@ 1 Introduction
${\mathbb{E}}[\prod_{i\in L}X_{i}]\leq\prod_{i\in L}{\mathbb{E}}[X_{i}]=\prod_{i\in L}x_{i}$ (1)
@@
One powerful setting for dependent rounding is the bipartite rounding algorithm of [5]. Namely, we have a bipartite graph $G=(U\cup V,E)$, where each edge $e$ has an associated weight $y_{e}\in[0,1]$. The goal is to generate random variables $Y\in\{0,1\}^{E}$ which match the fractional variables in expectation, while also satisfying certain nonpositive-correlation conditions. In the original algorithm of [5], which has found numerous applications in optimization problems, the algorithm guaranteed a limited case of Eq. (1): namely, it held for any edge set $L$ which was a subset of the neighborhood of any vertex.
For some combinatorial problems, nonpositive-correlation is not enough; we need to get strong negative correlation. Namely, for certain edge sets $L$, we want a stronger bound of the form
${\mathbb{E}}[\prod_{i\in L}X_{i}]\leq(1-\phi)\cdot\prod_{i\in L}x_{i}$ (2)
@@
One particularly interesting application comes from job scheduling on unrelated machines to minimize completion time. We will discuss this problem in more detail later, but for now let us provide a brief summary. We have a set of machines $\mathcal{M}$ and a set of jobs $\mathcal{J}$, where each job $j\in\mathcal{J}$ has a given processing time on each machine $i\in\mathcal{M}$. We want to minimize the objective function $\sum_{j\in\mathcal{J}}w_{j}C_{j}$, where $C_{j}$ is the completion time of $j$ on its chosen machine and $w_{j}$ is a weight function. A long series of approximation algorithms have been developed for this problem, often based on sophisticated convex relaxations and dependent-rounding algorithms; most recently, [9] achieved a $1.45$-approximation factor.
The aim of this paper is to obtain a more unified and general picture of bipartite dependent rounding with strong negative correlation. This algorithm, along with some other improvements, leads to a $1.398$-approximation algorithm for the above machine-scheduling problem, which improves over the algorithm of [9] while also being simpler and more generic.
@@ 1.1 Definitions and results for the bipartite-rounding setting
For combinatorial assignment problems, such as machine-scheduling, it can be useful to abstract to a setting of dependent rounding for bipartite graphs. Here, we are given a simple bipartite graph $G=(U\cup V,E)$. We call $U$ and $V$ left-nodes and right-nodes respectively. For any vertex $w\in U\cup V$ we define $\Gamma(w)$ to be the edges incident to $w$.
We suppose we are also given a fractional vector $\vec{x}\in[0,1]^{E}$ with $\sum_{e\in\Gamma(v)}x_{v}\leq 1$ for all right-nodes $v\in V$. Our goal is to generate rounded random variables $\vec{X}\in\{0,1\}^{E}$ which match the fractional variables, while also achieving nonpositive-correlation among certain subset of values $X_{e}$.
In Sections 2 and 3, we develop a general rounding algorithm based on negatively correlated Exponential random variables. These sections are completely self-contained and do not involve machine-scheduling in any way. We do this in two stages. First, we give a general method of generating Exponential random variables with certain types of negative correlation. Next, we use this in a simple contention-resolution scheme for the bipartite rounding, which takes as input a “rate” vector $\vec{\rho}\in[0,1]^{E}$ satisfying $\sum_{e\in\Gamma(u)}\rho_{e}\leq 1$ for all left-nodes $u$. This vector $\vec{\rho}$ is not necessarily related to $\vec{x}$; it can be carefully chosen to obtain different “shapes” of negative correlation.
For edges $e=(u,v),e^{\prime}=(u^{\prime},v^{\prime})$, let us define symmetric relation $e\sim e^{\prime}$ if (i) $u\neq u^{\prime},v^{\prime}\neq v$ and (ii) either $(u,v^{\prime})\in E$ or $(u^{\prime},v)\in E$. Equivalently, $e\sim e^{\prime}$ if $e,e^{\prime}$ have distance exactly two in the line graph of $G$. We say a set of edges $L\subseteq E$ is stable if there is no pair $e,e^{\prime}\in L$ with $e\sim e^{\prime}$.
@@
The algorithm DepRound satisfies the following rounding properties (A1) — (A4):
@@
    For any pair of edges $e_{1}=(u,v_{1}),e_{2}=(u,v_{2})$ which share a common left-node and which have $\rho_{e_{1}},\rho_{e_{2}}\in(0,1)$, we have ${\mathbb{E}}[X_{e_{1}}X_{e_{2}}]\leq(1-\phi_{e_{1},e_{2}})\cdot x_{e_{1}}x_{e_{2}}$, where we define the anti-correlation parameter $\phi_{e_{1},e_{2}}\in(0,1)$ by:
    $\phi_{e_{1},e_{2}}=\frac{((1-\rho_{e_{1}})^{1-1/x_{e_{1}}}-1)((1-\rho_{e_{2}})^{1-1/x_{e_{2}}}-1)}{(1-\rho_{e_{1}})^{1-1/x_{e_{1}}}(1-\rho_{e_{2}})^{1-1/x_{e_{2}}}+\rho_{e_{1}}+\rho_{e_{2}}-1}$
These guarantees are extremely general, but hard to interpret. We can make it more concrete in two ways. First, instead of considering stable edge sets, we can consider a setting where each left-node has some “blocks” of edges; edges within a block should satisfy strong negative correlation and edges incident to a left-node should satisfy nonpositive-correlation [1, 2, 8]. By transforming each block in the graph to a new separate left-node, we can get the following corollary:
@@
Let $G$ be a complete bipartite graph where each left-node $u\in U$ has a partition of its edges into blocks $B_{u,1},\dots,B_{u,\ell_{u}}$, with a given vector $\vec{\rho}\in[0,1]^{E}$ such that $\sum_{e\in B_{u,i}}\rho_{e}\leq 1$ for all left-nodes $u$ and blocks $i$. Then algorithm DepRound can be used to round the variables $\vec{x}$ into $\vec{X}\in\{0,1\}^{E}$ to satisfy the following properties (A1), (A2), (A3’), (A4’) as follows:
  • (A1)
  • (A2)
    For each edge $e$ there holds ${\mathbb{E}}[X_{e}]=x_{e}$
  • (A3’)
    For any left-node $u$ and edge-set $L\subseteq\Gamma(u)$, there holds ${\mathbb{E}}[\prod_{e\in L}X_{e}]\leq\prod_{e\in L}x_{e}$
  • (A4’)
    For any pair of edges $e_{1},e_{2}$ which are in some common block $B_{u,i}$ and which have $\rho_{e_{1}},\rho_{e_{2}}\in(0,1)$, we have ${\mathbb{E}}[X_{e_{1}}X_{e_{2}}]\leq(1-\phi_{e_{1},e_{2}})\cdot x_{e_{1}}x_{e_{2}}$, where $\phi_{e_{1},e_{2}}$ is defined as in Theorem 1.
There is also a question of choosing the parameter $\rho$. Intuitively, $\rho_{e}$ controls how much anti-correlation $e$ should have with other edges. We emphasize that there is no single optimal choice, and our scheduling algorithm will use this flexibility in a somewhat subtle way. However, there are a few natural parameterizations, as given in the following result:
Corollary 3.
Let $\phi_{e_{1},e_{2}}$ be the anti-correlation parameter of Theorem 1.
For a left-node $u$, let $\lambda_{u}=\sum_{e\in\Gamma(u)}x_{e}$. If we set $\rho_{e}=1-\mathrm{e}^{-x_{e}/\lambda_{u}}$ for all $e\in\Gamma(u)$, then [^3] Here and throughout we write $\mathrm{e}=2.718...$ Note that we use a slightly different font to distinguish the constant $\mathrm{e}$ from an edge $e$ in a graph.
$\phi_{e_{1},e_{2}}\geq\frac{1+\mathrm{e}^{1/\lambda_{u}}-\mathrm{e}^{x_{e_{1}}/\lambda_{u}}-\mathrm{e}^{x_{e_{2}}/\lambda_{u}}}{1+\mathrm{e}^{1/\lambda_{u}}}$
@@
$\phi_{e_{1},e_{2}}\geq\frac{\mathrm{e}^{1/\lambda_{u}}-1}{\mathrm{e}^{1/\lambda_{u}}+1}-0.57(x_{e_{1}}+x_{e_{2}})\max\{0,\lambda_{u}-0.45\}$
The first result strictly generalizes the result of [8] (which only considered the setting with $\lambda_{u}=1$). The second result can be much stronger quantitatively, and also has a simpler algebraic formula. We emphasize that the dependent rounding algorithm, and its analysis, is completely self-contained and does not depend on the machine-scheduling setting. Because of its generality, it may be applicable to other combinatorial optimization problems.
@@ 1.2 Machine-scheduling on unrelated machines
As shown in [1], going beyond this $1.5$-approximation ratio demands much more involved algorithms: some of the main convex relaxations have integrality gap $1.5$, and furthermore rounding strategies which treat each job independently, as had been used in all previous algorithms, are inherently limited to approximation ratio $1.5$. In a breakthrough, [1] achieved a $1.5-\varepsilon$ approximation factor for some minuscule constant $\varepsilon>0$. Since then, the approximation ratio has been further improved [11, 8, 9]; most recently, [9] gave a $1.45$-approximation (see also [12] for a faster implementation of that algorithm).
These newer algorithms can all be described in the same general framework. First, they solve an appropriate relaxation, giving fractional assignments $x^{(i)}_{j}$ (with perhaps some additional information). Second, they group the jobs on each machine, forming “clusters” of jobs with similar processing times. Finally, they apply some form of dependent rounding with strong negative correlation properties within each cluster to convert this into an integral assignment.
@@
First, the DepRound algorithm is much more flexible for rounding: its negative correlation guarantees scale with the size of each cluster or even the individual items, and are not determined by worst-case bounds on maximum cluster size. The parameter $\rho$ available in DepRound allows us to put less “anti-correlation strength” on the later jobs within a cluster. This makes the clusters act in a significantly more uniform way irrespective of their total mass. In particular, we no longer need to deal separately with jobs which have “large” mass on a given machine.
@@ 2 Negatively-correlated Exponential random variables
Before we consider bipartite rounding, we derive a more basic result in probability theory: how to generate Exponential random variables with strong negative correlations. The analysis depends heavily on properties of negatively-associated (NA) random variables [10]. Formally, we say that random variables $X_{1},\dots,X_{k}$ are NA if for any disjoint subsets $A_{1},A_{2}\subseteq\{1,...,k\}$ and any increasing functions $f_{1},f_{2}$, there holds
${\mathbb{E}}[f_{1}(X_{i}:i\in A_{1})f_{2}(X_{j}:j\in A_{2})]\leq{\mathbb{E}}[f_{1}(X_{i}:i\in A_{1})]{\mathbb{E}}[f_{2}(X_{j}:j\in A_{2})]$ (3)
@@
1 Set $X\leftarrow\textsc{MultivariateGeometric}(\vec{\rho})$.
2 for $i\in[n]$ do
@@ 6             Set $Z_{i}=\alpha_{i}(X_{i}+S_{i})$
9             Draw $Z_{i}$ as an independent Exponential random variable with rate $1$
10
Return $Z$
[…]
The algorithm can be implemented to run in randomized polynomial time via standard sampling procedures. We now analyze its probabilistic properties.
@@ 9             Draw $Z_{i}$ as an independent Exponential random variable with rate $1$
It is convenient to add a dummy element 00 with $\rho_{0}=1-\sum_{i=1}^{n}\rho_{i}$. We first claim that random variables $X_{1},\dots,X_{n}$ are NA. To see this, observe that $X_{1},\dots,X_{n}$ can be viewed in terms of the following infinite sequence of random variables: for each $j\in\mathbb{Z}_{\geq 0}$, draw variable $F_{j}\in\{0,\dots,n\}$ with $\Pr(F_{j}=i)=\rho_{i}$ for all $i,j$. Then set $X_{i}=\min\{j:F_{j}=i\}$ for each $i$.
Now let $B_{i,j}$ be the indicator random variable for the event $F_{j}=i$. For each $j$, the values $B_{i,j}$ are zero-one random variables which sum to one. Hence, for fixed $j$ they are NA. Since there is no interaction between the different indices $j$, all random variables $B_{i,j}$ are NA. Furthermore, each $X_{i}$ is a monotone-down function of the values $B_{i,j}$. Hence, all variables $X_{i}$ are NA.
The $S_{i}$ variables are clearly NA since they are independent. Each variable $Z_{i}$ for $\rho_{i}\in(0,1)$ is an increasing function of the random variables $S_{i},X_{i}$. Hence, these $Z$ variables are also NA. The $Z_{i}$ variables with $\rho_{i}\in\{0,1\}$ are NA since they independent of all other variables. ∎
@@
${\mathbb{E}}[\mathrm{e}^{q_{1}Z_{1}+q_{2}Z_{2}}]={\mathbb{E}}[\mathrm{e}^{\theta_{1}L_{1}+\theta_{2}L_{2}}]={\mathbb{E}}[\mathrm{e}^{\theta_{1}S_{1}+\theta_{2}S_{2}}\cdot\mathrm{e}^{\theta_{1}X_{1}+\theta_{2}X_{2}}]={\mathbb{E}}[\mathrm{e}^{\theta_{1}S_{1}}]{\mathbb{E}}[\mathrm{e}^{\theta_{2}S_{2}}]{\mathbb{E}}[\mathrm{e}^{\theta_{1}X_{1}+\theta_{2}X_{2}}]$
@@
[…]

### v8

@@ Abstract
We describe a new dependent-rounding algorithmic framework for bipartite graphs. Given a fractional assignment $\vec{x}$ of values to edges of a graph $G=(U\cup V,E)$, the algorithms return an integral solution $\vec{X}$ such that each right-node $v\in V$ has at most one neighboring edge $f$ with $X_{f}=1$, and the variables $X_{e}$ also satisfy broad nonpositive-correlation properties. In particular, for any edges $e_{1},e_{2}$ sharing a left-node $u\in U$, the variables $X_{e_{1}},X_{e_{2}}$ have strong negative correlation, i.e. the expectation of $X_{e_{1}}X_{e_{2}}$ is significantly below $x_{e_{1}}x_{e_{2}}$.
@@ 1 Introduction
${\mathbb{E}}\Bigl[\prod_{i\in L}X_{i}\Bigr]\leq\prod_{i\in L}{\mathbb{E}}[X_{i}]=\prod_{i\in L}x_{i}$ (1)
@@
For some combinatorial problems, nonpositive-correlation is not enough; we need strong negative correlation. Namely, for certain edge sets $L$, we want a bound of the form
${\mathbb{E}}\Bigl[\prod_{i\in L}X_{i}\Bigr]\leq(1-\phi)\cdot\prod_{i\in L}x_{i}$ (2)
@@
One particularly interesting application comes from job scheduling on unrelated machines to minimize completion time. We will discuss this problem in more detail later, but for now let us provide a brief summary. We have a set of machines $\mathcal{M}$ and a set of jobs $\mathcal{J}$, where each job $j\in\mathcal{J}$ has a given processing time on each machine $i\in\mathcal{M}$. We want to minimize the objective function $\sum_{j\in\mathcal{J}}w_{j}C_{j}$, where $C_{j}$ is the completion time of $j$ on its chosen machine and $w_{j}$ is a weight function. A long series of approximation algorithms have been developed for this problem, often based on sophisticated convex relaxations and dependent-rounding algorithms.
The aim of this paper is to obtain a more unified and general picture of dependent rounding with strong negative correlation. Our algorithm, along with some other improvements, leads to a $1.398$-approximation algorithm for the above machine-scheduling problem, which improves over the algorithm of [10] while also being simpler and more generic.
@@ 1.1 Definitions and results for the bipartite-rounding setting
Many combinatorial assignment problems, such as machine-scheduling, can be abstracted in the following framework: we consider a bipartite graph $G=(U\cup V,E)$, where the fractional variables $\vec{x}\in[0,1]^{E}$ are associated to its edges. Our goal is to generate rounded random variables $\vec{X}\in\{0,1\}^{E}$ which match the fractional variables, while also achieving nonpositive-correlation or strong negative correlation among certain subset of values $X_{e}$ corresponding to the vertices.
We call $U$ and $V$ left-nodes and right-nodes respectively. For any vertex $w\in U\cup V$ we define $\Gamma(w)$ to be the edges incident to $w$.
The bipartite rounding algorithm of [6] is a powerful example of this framework. This algorithm, which has found numerous applications in optimization problems, guaranteed a limited case of Eq. (1): namely, it held for any edge set $L$ which was a subset of the neighborhood of any vertex.
Our first main contribution, in Sections 2 and 3, is to develop a new bipartite-graph rounding algorithm based on negatively correlated Exponential random variables. The construction has two stages. First, we give a general method of generating Exponential random variables with certain types of negative correlation. Second, we use this in a graph contention-resolution scheme, which takes as input a “rate” vector $\vec{\rho}\in[0,1]^{E}$; we will discuss its role shortly.
Let us say that a set of edges $L\subseteq E$ is stable if it does not contain any pair of edges of the form $e=(u,v),e^{\prime}=(u^{\prime},v^{\prime})$, where $(u,v^{\prime})$ is also an edge in $E$. Equivalently, $L$ does not contain any pair of edges whose distance is precisely two in the line graph of $G$.
@@
Suppose $\sum_{e\in\Gamma(v)}x_{v}\leq 1$ for all right-nodes $v\in V$, and $\sum_{e\in\Gamma(u)}\rho_{u}\leq 1$ for all left-nodes $u\in U$. Then the algorithm DepRound satisfies the following properties (A1) — (A4):
@@
    For any pair of edges $e_{1}=(u,v_{1}),e_{2}=(u,v_{2})$ sharing a common left-node, we have
    ${\mathbb{E}}[X_{e_{1}}X_{e_{2}}]\leq(1-\Phi(x_{e_{1}},x_{e_{2}};\rho_{e_{1}},\rho_{e_{2}}))x_{e_{1}}x_{e_{2}},$
    where we define the function $\Phi(x_{1},x_{2};\rho_{1},\rho_{2}):[0,1]^{4}\rightarrow[0,1]$ by
    $\Phi(x_{1},x_{2};\rho_{1},\rho_{2})=\begin{cases}\frac{((1-\rho_{1})^{1-1/x_{1}}-1)((1-\rho_{2})^{1-1/x_{2}}-1)}{(1-\rho_{1})^{1-1/x_{1}}(1-\rho_{2})^{1-1/x_{2}}+\rho_{1}+\rho_{2}-1}&\text{for $x_{1},x_{2},\rho_{1},\rho_{2}\in(0,1)$}\\ 0&\text{otherwise}\end{cases}$
There is the question of choosing the parameter $\rho$. This vector $\vec{\rho}$ is not necessarily related to $\vec{x}$; it can be carefully chosen to obtain different “shapes” of negative correlation. Intuitively, $\rho_{e}$ controls how much anti-correlation $e$ should have with other edges. We emphasize that there is no single optimal choice, and our scheduling algorithm will use this flexibility in a somewhat subtle way. However, there are a few natural parameterizations, as given in the following result:
@@
For a left-node $u$, let $\lambda_{u}=\sum_{e\in\Gamma(u)}x_{e}$. If we set $\rho_{e}=1-\mathrm{e}^{-x_{e}/\lambda_{u}}$ for all $e\in\Gamma(u)$, then for any edges $e_{1},e_{2}$ with $x_{e_{1}}x_{e_{2}}>0$, we have [^3] Here and throughout we write $\mathrm{e}=2.718$... We use a different font to distinguish the constant $\mathrm{e}$ from an edge $e$ in a graph.
$\Phi(x_{e_{1}},x_{e_{2}};\rho_{e_{1}},\rho_{e_{2}})\geq 1-\frac{\mathrm{e}^{x_{e_{1}}/\lambda_{u}}+\mathrm{e}^{x_{e_{2}}/\lambda_{u}}}{1+\mathrm{e}^{1/\lambda_{u}}}$
@@
$\Phi(x_{e_{1}},x_{e_{2}};\rho_{e_{1}},\rho_{e_{2}})\geq\frac{\mathrm{e}^{1/\lambda_{u}}-1}{\mathrm{e}^{1/\lambda_{u}}+1}-0.57(x_{e_{1}}+x_{e_{2}})\max\{0,\lambda_{u}-0.45\}$
The first result strictly generalizes the result of [9] (which only considered the setting with $\lambda_{u}=1$). The second result can be stronger quantitatively, and also has a simpler algebraic formula.
We note that there is a closely related setting, as originally considered in [2, 3, 9], where we have a bipartite graph $G$ where the edges incident to each left-node are partitioned into “blocks.” In this setting, we should have nonpositive-correlation among edges incident to each left-node, and strong negative correlation within each block. Our framework can also handle this setting, by transforming each block of the graph $G$ to a new separate left-node. Note that the neighborhood of a vertex in the original graph $G$ will form a stable edge-set in the resulting graph $G^{\prime}$.
We emphasize that the dependent rounding algorithm, and its analysis, is completely self-contained and does not depend on the machine-scheduling setting. Because of its generality, it may be applicable to other combinatorial optimization problems.
@@ 1.2 Machine-scheduling on unrelated machines
As shown in [2], going beyond this approximation ratio demands much more involved algorithms: some of the main convex relaxations have integrality gap $1.5$, and furthermore rounding strategies which treat each job independently, as had been used in all previous algorithms, are inherently limited to approximation ratio $1.5$. In a breakthrough, [2] achieved a $1.5-\varepsilon$ approximation factor for some minuscule constant $\varepsilon>0$. There have been further improvements [12, 9, 10]; most recently before this work, [10] gave a $1.45$-approximation (see also [13] for a faster implementation of that algorithm).
These newer algorithms can all be described in the same general framework. First, they solve an appropriate relaxation, giving fractional assignments $x^{(i)}_{j}$. Second, they group the jobs on each machine, forming “clusters” of jobs with similar processing times. Finally, they apply some form of dependent rounding with strong negative correlation properties within each cluster to convert this into an integral assignment.
@@
First, the DepRound algorithm is much more flexible for rounding: its negative correlation guarantees scale with the size of each cluster or even the individual items, as opposed to the maximum cluster size. The parameter $\rho$ available in DepRound allows us to put less “anti-correlation strength” on the final job within a cluster. This makes the clusters act in a significantly more uniform way irrespective of their total mass. In particular, we no longer need to deal separately with jobs which have “large” mass on a given machine.
@@
Note:
After the conference version of this paper, [14] proposed a new algorithm for this problem with approximation ratio $1.36$. It uses different techniques, based on iterated rounding. Its approximation ratio (and integrality gap) are shown in terms of the Configuration LP, which is stronger than either the time-indexed LP or SDP relaxations.
@@ 2 Negatively-correlated Exponential random variables
Before we consider bipartite rounding, we derive a more basic result in probability theory: how to generate Exponential random variables with strong negative correlations. The analysis relies heavily on properties of negatively-associated (NA) random variables [11]. Formally, we say that random variables $X_{1},\dots,X_{k}$ are NA if for any disjoint subsets $A,B\subseteq\{1,...,k\}$ and any increasing functions $f,g$, there holds
${\mathbb{E}}[f(X_{i}:i\in A)g(X_{j}:j\in B)]\leq{\mathbb{E}}[f(X_{i}:i\in A)]{\mathbb{E}}[g(X_{j}:j\in B)]$ (3)
@@
Our algorithm also makes use of the Multivariate Geometric distribution [1]: given a probability vector $\vec{p}\in[0,1]^{n}$, this distribution on vector $(X_{1},\dots,X_{n})$ is defined by running an infinite sequence of experiments, where each $j^{\text{th}}$ experiment takes on value $i$ with probability $p_{i}$, and setting $X_{i}$ to be the number of trials before first seeing $i$. We note a few simple properties of this distribution.
Observation 5.
Let $(X_{1},\dots,X_{n})$ be a Multivariate Geometric random variable with rate $\vec{p}$. Then the variables $X_{1},\dots,X_{n}$ are NA and each marginal distribution $X_{i}$ is Geometric with rate $p_{i}$.
Proof.
Observe that $X_{1},\dots,X_{n}$ can be viewed in terms of the following infinite sequence of random variables: for each $j\in\mathbb{Z}_{\geq 0}$, draw variable $F_{j}\in\{1,\dots,n\}$ with $\Pr(F_{j}=i)=p_{i}$ for all $i,j$. Then set $X_{i}=\min\{j:F_{j}=i\}$ for each $i$. Let $B_{i,j}$ be the indicator random variable for the event $F_{j}=i$. For each $j$, the values $B_{i,j}$ are zero-one random variables which sum to one, so they are NA. Since there is no interaction between the different indices $j$, all random variables $B_{i,j}$ are NA. Furthermore, each $X_{i}$ is a monotone-down function of the values $B_{i,j}$. Hence, all variables $X_{i}$ are NA.
Furthermore, each $X_{i}$ is the waiting time until we see the first variable $B_{i,j}$; since these are Bernoulli with rate $p_{i}$, the distribution of $X_{i}$ is Geometric with rate $p_{i}$
@@
1Set $\rho_{0}=1-(\rho_{1}+\dots+\rho_{n})$
2 Set $(X_{0},\dots,X_{n})\leftarrow\textsc{MultivariateGeometric}(\vec{\rho})$.
3 for $i=1,\dots,n$ do
@@ 7      Set $Z_{i}=\alpha_{i}(X_{i}+S_{i})$
10      Draw $Z_{i}$ as an independent Exponential random variable with rate 11
11
return vector $Z$
Algorithm 1 $\textsc{CorrelatedExponential}(\vec{\rho})$
The density function at Line 5 is valid in that $\int_{s=0}^{1}\alpha_{i}\mathrm{e}^{-\alpha_{i}s}/\rho_{i}\ \mathrm{d}s=\frac{1}{\rho_{i}}(1-\mathrm{e}^{-\alpha_{i}})=1$. The algorithm can be implemented to run in randomized polynomial time via standard sampling procedures. We now analyze its probabilistic properties.
@@ 10      Draw $Z_{i}$ as an independent Exponential random variable with rate 11
The random variables $X_{i}$ are NA by Observation 5. The $S_{i}$ variables are clearly NA since they are independent. Each variable $Z_{i}$ for $\rho_{i}\in(0,1)$ is an increasing function of the random variables $S_{i},X_{i}$. Hence, these $Z$ variables are also NA. The $Z_{i}$ variables with $\rho_{i}\in\{0,1\}$ are NA since they independent of all other variables. ∎
@@
${\mathbb{E}}[\mathrm{e}^{q_{1}Z_{1}+q_{2}Z_{2}}]={\mathbb{E}}[\mathrm{e}^{\theta_{1}L_{1}+\theta_{2}L_{2}}]={\mathbb{E}}[\mathrm{e}^{\theta_{1}S_{1}}]{\mathbb{E}}[\mathrm{e}^{\theta_{2}S_{2}}]{\mathbb{E}}[\mathrm{e}^{\theta_{1}X_{1}+\theta_{2}X_{2}}]$
@@
${\mathbb{E}}[\mathrm{e}^{\theta_{i}S_{i}}]$ $=\frac{\alpha_{i}}{\rho_{i}}\int_{s=0}^{1}\mathrm{e}^{-\alpha_{i}s+\theta_{i}s}\ \mathrm{d}s=\frac{\alpha_{i}(1-\mathrm{e}^{\theta_{i}-\alpha_{i}})}{\rho_{i}(\alpha_{i}-\theta_{i})}$